# Generalised Bayesian Inference using Robust divergences for von Mises-Fisher distribution

### v1

@@ Abstract
This paper focusses on robust estimation of location and concentration parameters of the von Mises-Fisher distribution in the Bayesian framework. The von Mises-Fisher (or Langevin) distribution has played a central role in directional statistics. Directional data have been investigated for many decades, and more recently, they have gained increasing attention in diverse areas such as bioinformatics and text data analysis. Although outliers can significantly affect the estimation results even for directional data, the treatment of outliers remains an unresolved and challenging problem. In the frequentist framework, numerous studies have developed robust estimation methods for directional data with outliers, but, in contrast, only a few robust estimation methods have been proposed in the Bayesian framework. In this paper, we propose Bayesian inference based on density power-divergence and $\gamma$-divergence and establish their asymptotic properties and robustness. In addition, the Bayesian approach naturally provides a way to assess estimation uncertainty through the posterior distribution, which is particularly useful for small samples. Furthermore, to carry out the posterior computation, we develop the posterior computation algorithm based on the weighted Bayesian bootstrap for estimating parameters. The effectiveness of the proposed methods is demonstrated through simulation studies. Using two real datasets, we further show that the proposed method provides reliable and robust estimation even in the presence of outliers or data contamination.
@@ 1 Introduction
$f(\bm{x}\mid\bm{\mu},\kappa)=\frac{\kappa^{(p-2)/2}}{(2\pi)^{p/2}I_{(p-2)/2}(\kappa)}\exp\{\kappa\bm{\mu}^{\top}\bm{x}\},\quad\bm{x}\in\mathcal{S}_{p},$ (1)
@@
In the Bayesian context, the location and concentration parameters of the von Mises-Fisher distribution have also long been studied in the literature (see, for example, Damien and Walker,, 1999; Giummolè et al.,, 2019; Hornik and Grün,, 2013; Mardia and El-Atoum,, 1976; Pal and Gaskins,, 2022). However, robust estimation for directional data has not been studied enough, and the existing literature in Bayesian inference is very limited. In Bayesian statistics, robustness against outliers is also an important issue (see, e.g., Berger et al.,, 1994; O’Hagan and Pericchi,, 2012, among others). Over the past few years, a number of robust Bayesian estimation methods have been proposed, including approaches based on heavy-tailed distributions to accommodate outliers (e.g., Desgagné,, 2015; Gagnon et al.,, 2020; Hamura et al.,, 2022), divergence-based methods (e.g., Futami et al.,, 2018; Ghosh and Basu,, 2016; Hooker and Vidyashankar,, 2014; Jewson et al.,, 2018; Momozaki and Nakagawa,, 2023; Nakagawa and Hashimoto,, 2020, 2021), and loss-based approaches (e.g., Kawakami and Hashimoto,, 2023; Matsubara et al.,, 2022, among others). Nevertheless, these studies have primarily focused on data in Euclidean space, and have not considered data defined on compact sets such as directional data. Motivated by this perspective, we focus on divergence-based generalized Bayesian inference for directional data and consider robust loss functions derived from the density power divergence and the $\gamma$-divergence.
@@
This paper provides a robust Bayesian inference framework for directional data using divergences. In this study, we propose the generalised posterior distribution for directional data using density power-divergence and $\gamma$-divergence and investigate their robustness. This approach facilitates a straightforward assessment of estimation uncertainty, particularly in scenarios characterized by a limited sample size. Additionally, we provide the posterior computation algorithm based on the Bayesian bootstrap (Lyddon et al.,, 2019; Newton and Raftery,, 1994; Newton et al.,, 2021; Rubin,, 1981) to sample from the proposed posterior distribution. This makes it possible to approximately sample from the posterior distribution even when $p$ is large, thereby making Bayesian inference feasible. This paper is organized as follows: In Section 2, we explain the von Mises-Fisher distribution and its properties, which were used as a model for the directional data in the posterior distribution of this study. In Section 3, we propose a robust posterior distribution based on the density power-divergence and the $\gamma$-divergence. We also discuss the method for computing robust posterior distributions through sampling. While sampling methods based on MCMC are typically employed, they were not applicable to the current model due to specific application constraints. Therefore, parameter computations were carried out using the Bayesian Bootstrap method in this study. Section 4 discuss a robustness of the proposed estimators using the robust posterior distributions in directional data. We derive the standardized influence function of the proposed estimators and compare them with ordinary Bayesian inference. In Section 5, we present that numerical performance of the posterior means based on the density power-divergence and $\gamma$-divergence in simulation studies, and we see that these posterior means are hardly influenced even in presence of outliers. In Section 6, we apply our proposed estimation to the wind direction data and the gene expression data, illustrating its effectiveness in practical settings.
2 Robust estimation for von Mises-Fisher distribution
@@
It is well known that the MLE is highly sensitive to outliers. This has led to the development of many robust estimation methods to address this issue (see, for example, Hampel et al.,, 1986; Huber and Ronchetti,, 2009; Maronna et al.,, 2019). One well-known robust estimation method is the divergence approach. For the probability density functions $f$ and $g$, Basu et al., (1998) proposed the density power divergence defined as $\mathcal{D}_{\alpha}(g,f)=-d_{\alpha}(g,g)+d_{\alpha}(g,f)$ where $d_{\alpha}(g,f)$ is the the cross entropy given as
@@
Jones et al., (2001) also proposed the $\gamma$-divergence defined as $\mathcal{D}_{\gamma}(g,f)=-d_{\gamma}(g,g)+d_{\gamma}(g,f)$ where $d_{\alpha}(g,f)$ is the the cross entropy given as
$d_{\gamma}(g,f)$ $=-\frac{1}{\gamma}\log\int g(\bm{x})f(\bm{x})^{\gamma}{\rm d}\bm{x}+\frac{1}{(1+\gamma)}\log\int f\bm{x})^{1+\gamma}{\rm d}\bm{x},\quad(\gamma>0).$
Here, these are consistent with Kullback-Leibler (KL) divergence when $\alpha,\gamma=0$. Basu et al., (1998) and Jones et al., (2001) proposed estimators by minimizing the empirical versions of these divergences and their estimators are known to hold strong robustness.
@@
It is noted that $\bm{\xi}$ is a parameter in $\mathbb{R}^{p}$, and the Euclidean norm of $\bm{\xi}$ represents the concentration parameter $\kappa$, while the standardized parameter $\bm{\xi}/\|\bm{\xi}\|$ represents the mean direction parameter $\bm{\mu}$. Under this parametrisation, the cross entropy of the density power divergence between the von Mises-Fisher distribution $f_{\bm{\xi}}$ and an arbitrary density $g$ on $\mathcal{S}_{p}$ is given as follows
@@
where $K_{p}(\bm{\xi})=\{(2\pi)^{p/2}I_{(p-2)/2}(\|\bm{\xi}\|)\}/\|\bm{\xi}\|^{(p-2)/2}$. Similarly, the cross entropy of $\gamma$-divergence between $f_{\bm{\xi}}$ and an arbitrary density $g$ is given by
@@
3 Robust Generalised Bayesian Inference for von Mises-Fisher distribution
Based on Bissiri et al., (2016); Bissiri and Walker, (2019), the general Bayesian framework achieves belief updating while satisfying coherence and additivity for parameters of interest in the $M$-estimation. Then the generalised posterior distribution is given as a functional form for belief updating. This implies that the generalised posterior provides a probabilistic quantification of the uncertainty associated with the $M$-estimator. Consequently, we apply generalised Bayesian inference to the von Mises-Fisher distribution using robust divergences.
3.1 Generalized Posterior distributions base on robust divergences
@@ 3.2 Asymptotic Properties
as $n\rightarrow\infty$, where $\phi(\cdot;A)$ is the density function of a p-variate normal distribution with zero mean vector and covariance matrix $A$, and $\pi^{*(d)}(\bm{t}_{n}\mid\bm{x}_{1:n})$ is the generalized posterior density function of the normalized random variable $\bm{t}_{n}=(t_{1},\ldots,t_{p})^{\top}=\sqrt{n}(\bm{\xi}-\hat{\bm{\xi}}_{n}^{(d)})$ given $\bm{x}_{1:n}$.
@@ 3.3 Posterior Computation
In this section, we consider approximate posterior sampling by optimising randomly weighted likelihood functions using the weighted Bayesian bootstrap (WBB) method. Generally, sampling from the posterior distribution is typically performed using MCMC methods, such as Hamiltonian Monte Carlo, the Metropolis-Hastings algorithm, and Gibbs sampling. In directional statistics, numerous studies have developed MCMC for the von Mises-Fisher distribution (Nuñez-antonio and Gutiérrez-peña,, 2005; Pal and Gaskins,, 2022), and a general approach to MCMC simulation on embedded Riemannian manifolds was introduced by Byrne and Girolami, (2013); Habeck et al., (2023). However, implementing these MCMC methods for the robust divergence-based generalised posterior distributions considered in this study is extremely difficult, because the resulting generalised posterior involves special functions and therefore has a highly complicated form. Although tools such as Stan (Carpenter et al.,, 2017) have made MCMC much easier to use in recent years, they cannot be applied in our setting: the DPD and $\gamma$-D posteriors for the von Mises-Fisher distribution involve Bessel functions of order (p2)/2(p-2)/2, and Stan’s built-in Bessel functions do not support fractional orders. Therefore, Stan cannot be used when the dimension $p$ is odd.
@@ 4 Robustness
Under the quadratic loss function, the DPD Bayes functional $T_{n}^{(\alpha)L}(G)$ and the $\gamma$-D Bayes functional $T_{n}^{(\gamma)L}(G)$ are the DPD posterior mean functional $T_{n}^{(\alpha)e}(G)$ and the $\gamma$-posterior mean functional $T_{n}^{(\gamma)e}(G)$, respectively. $T_{n}^{(\alpha)e}(G)$ and $T_{n}^{(\gamma)e}(G)$ are given as
@@ 6.1 Wind direction data
We used the wind dataset provided in the circular R package (Agostinelli and Lund,, 2025) to illustrate the basic characteristics of directional data. Figure 6 illustrates the posterior mean directions of $\bm{\mu}=(\mu_{1},\mu_{2})^{\top}$ estimated from real data using three different posteriors: the ordinary posterior (KL), the DPD posterior, and the $\gamma$-D posterior. Each arrow represents the estimated posterior mean direction on the unit circle, while the surrounding points indicate the sample observations. The tuning parameters $\alpha$ and $\gamma$ are chosen so that ARE is $95\%$, see Appendix A.3 for details.
From Figure 6, we can see that the ordinary posterior mean (KL; $0.29$ radian) is noticeably shifted due to the influence of outliers, whereas the DPD posterior mean (DPD; $0.17$ radian) and $\gamma$-D posterior mean (Gamma; $0.16$ radian) remain stable and close to the dominant direction of the data, demonstrating their robustness. The circular dispersion of the posterior samples for $\bm{\mu}$ was smallest for the DPD and $\gamma$-D posteriors, with mean resultant lengths of $0.9993$ and $0.9993$, compared to $0.9989$ for the KL posterior. Consistently, the estimated concentration parameters $\kappa$ were higher for the DPD and $\gamma$-D posteriors ($2.99$ and $3.69$, respectively) than for the ordinary posterior ($1.78$), confirming their robustness to outliers.
@@ 6.2 Gene expression data
Directional data also arise naturally in bioinformatics (e.g. Eisen et al.,, 1998), particularly in the analysis of gene expression data, in which the von Mises model has been applied for modeling directional structures (e.g. Banerjee et al.,, 2005). We used yeast gene expression data from the NCBI Gene Expression Omnibus (accession GSE12345) (Edgar et al.,, 2002), retrieved via the R package GEOquery (Davis and Meltzer,, 2007), and processed using Biobase from the Bioconductor project (Gentleman et al.,, 2004). To interpret gene expression profiles as directional data, each observation $\bm{z}_{i}=(z_{i1},\ldots,z_{ip})^{\top}$ is normalized to a vector $\bm{x}_{i}=(x_{i1},\ldots,x_{ip})^{\top}$ on the unit sphere by
@@
where $\bar{z}=p^{-1}\sum_{k=1}^{p}z_{ik}$. This normalization ensures that $|\bm{x}_{i}|_{2}=1$, so that each sample lies on the unit sphere. Under this transformation, the Pearson correlation between $\bm{z}_{i}$ and $\bm{z}_{j})$ can be equivalent to the cosine similarity between $\bm{x}_{i}$ and $\bm{x}_{j}$.
@@
Table 1 presents the angular distance and cosine distance between posterior mean directions obtain from the ordinary, DPD, and $\gamma$-D posteriors ($\alpha=\gamma=0.15$). Figure 8 also provides histograms of the cosine similarities between the data and these posterior mean directions. From the table and figure, we can see that the posterior mean directions are close to one another, suggesting minimal differences in the resulting point estimates. On the other hand, Table 2, which summarizes the posterior means of $\kappa$, shows that the ordinary posterior markedly underestimates $\kappa$ due to the strong influence of outliers. By contrast, the DPD and $\gamma$-D posteriors yield considerably more robust estimates.

### v2

# Divergence-based Robust Generalised Bayesian Inference for Directional Data via von Mises-Fisher models

@@ Abstract
This paper focusses on robust estimation of location and concentration parameters of the von Mises-Fisher distribution in the Bayesian framework. The von Mises-Fisher (or Langevin) distribution has played a central role in directional statistics. Directional data have been investigated for many decades, and more recently, they have gained increasing attention in diverse areas such as bioinformatics and text data analysis. Although outliers can significantly affect the estimation results even for directional data, the treatment of outliers remains an unresolved and challenging problem. In the frequentist framework, numerous studies have developed robust estimation methods for directional data with outliers, but, in contrast, only a few robust estimation methods have been proposed in the Bayesian framework. In this paper, we propose Bayesian inference based on the density power divergence and the $\gamma$-divergence and establish their asymptotic properties and robustness. In addition, the Bayesian approach naturally provides a way to assess estimation uncertainty through the posterior distribution, which is particularly useful for small samples. Furthermore, to carry out the posterior computation, we develop the posterior computation algorithm based on the weighted Bayesian bootstrap for estimating parameters. The effectiveness of the proposed methods is demonstrated through simulation studies. Using two real datasets, we further show that the proposed method provides reliable and robust estimation even in the presence of outliers or data contamination.
@@ 1 Introduction
$f(\bm{x}\mid\bm{\mu},\kappa)=\frac{\kappa^{(p-2)/2}}{(2\pi)^{p/2}I_{(p-2)/2}(\kappa)}\exp\{\kappa\bm{\mu}^{\top}\bm{x}\},\quad\bm{x}\in\mathcal{S}_{p},$
@@
In the Bayesian context, the location and concentration parameters of the von Mises-Fisher distribution have also long been studied in the literature (see, for example, Damien and Walker,, 1999; Giummolè et al.,, 2019; Hornik and Grün,, 2013; Mardia and El-Atoum,, 1976; Pal and Gaskins,, 2022). However, robust estimation for directional data has not been studied enough, and the existing literature in Bayesian inference is very limited. In Bayesian statistics, robustness against outliers is also an important issue (see, e.g., Berger et al.,, 1994; O’Hagan and Pericchi,, 2012, among others). Over the past few years, a number of robust Bayesian estimation methods have been proposed, including approaches based on heavy-tailed distributions to accommodate outliers (e.g., Desgagné,, 2015; Gagnon et al.,, 2020; Hamura et al.,, 2022), divergence-based methods (e.g., Futami et al.,, 2018; Ghosh and Basu,, 2016; Hooker and Vidyashankar,, 2014; Jewson et al.,, 2018; Momozaki and Nakagawa,, 2023; Nakagawa and Hashimoto,, 2020, 2021), and loss-based approaches (e.g., Kawakami and Hashimoto,, 2023; Matsubara et al.,, 2022, among others). Nevertheless, these studies have primarily focused on data in Euclidean space, and have not considered data defined on compact sets such as directional data. Motivated by this perspective, we focus on divergence-based generalised Bayesian inference for directional data and consider robust loss functions derived from the density power divergence and the $\gamma$-divergence.
@@
This paper provides a robust Bayesian inference framework for directional data using divergences via von Mises-Fisher models. In this study, we propose the generalised posterior distribution for directional data using the density power divergence and the $\gamma$-divergence and investigate their robustness. This approach facilitates a straightforward assessment of estimation uncertainty, particularly in scenarios characterized by a limited sample size. Additionally, we provide the posterior computation algorithm based on the Bayesian bootstrap (Lyddon et al.,, 2019; Newton and Raftery,, 1994; Newton et al.,, 2021; Rubin,, 1981) to sample from the proposed posterior distribution. This makes it possible to approximately sample from the posterior distribution even when $p$ is large, thereby making Bayesian inference feasible. This paper is organised as follows: In Section 2, we explain a robust estimation for the von Mises-Fisher models based on the density power divergence and the $\gamma$-divergence. In Section 3, we propose a robust posterior distribution based on the density power divergence and the $\gamma$-divergence. We also discuss the method for computing robust posterior distributions through sampling. While sampling methods based on MCMC are typically employed, they were not applicable to the proposed posterior distributions due to specific application constraints. Therefore, posterior computations were carried out using the Bayesian Bootstrap method in this study. Section 4 discusses robustness of the proposed estimators using robust posterior distributions in directional data. We derive the standardised influence function of the proposed estimators and compare them with ordinary Bayesian inference. In Section 5, we present the numerical performance of posterior means based on the density power divergence and the $\gamma$-divergence in simulation studies, and we see that these posterior means are hardly influenced even in the presence of outliers. In Section 6, we apply our proposed estimation to the wind direction data and the gene expression data, illustrating its effectiveness in practical settings.
2 Robust estimation for von Mises-Fisher models
@@
It is well known that the MLE is highly sensitive to outliers. This has led to the development of many robust estimation methods to address this issue (see, for example, Hampel et al.,, 1986; Huber and Ronchetti,, 2009; Maronna et al.,, 2019). One well-known robust estimation method is the divergence approach. For the probability density functions $f$ and $g$, Basu et al., (1998) proposed the density power divergence defined as $\mathcal{D}_{\alpha}(g,f)=-d_{\alpha}(g,g)+d_{\alpha}(g,f)$ where $d_{\alpha}(g,f)$ is the cross entropy given as
@@
Jones et al., (2001) also proposed the $\gamma$-divergence defined as $\mathcal{D}_{\gamma}(g,f)=-d_{\gamma}(g,g)+d_{\gamma}(g,f)$ where $d_{\alpha}(g,f)$ is the cross entropy given as
$d_{\gamma}(g,f)$ $=-\frac{1}{\gamma}\log\int g(\bm{x})f(\bm{x})^{\gamma}{\rm d}\bm{x}+\frac{1}{(1+\gamma)}\log\int f(\bm{x})^{1+\gamma}{\rm d}\bm{x},\quad(\gamma>0).$
Here, these are consistent with Kullback-Leibler (KL) divergence when $\alpha,\gamma=0$. Basu et al., (1998) and Jones et al., (2001) proposed estimators by minimising the empirical versions of these divergences and their estimators are known to hold strong robustness.
@@
It is noted that $\bm{\xi}$ is a parameter in $\mathbb{R}^{p}$, and the Euclidean norm of $\bm{\xi}$ represents the concentration parameter $\kappa$, while the standardised parameter $\bm{\xi}/\|\bm{\xi}\|$ represents the mean direction parameter $\bm{\mu}$. Under this parametrisation, the cross entropy of the density power divergence between the von Mises-Fisher distribution $f_{\bm{\xi}}$ and an arbitrary density $g$ on $\mathcal{S}_{p}$ is given as follows
@@
where $K_{p}(\bm{\xi})=\{(2\pi)^{p/2}I_{(p-2)/2}(\|\bm{\xi}\|)\}/\|\bm{\xi}\|^{(p-2)/2}$. Similarly, the cross entropy of the $\gamma$-divergence between $f_{\bm{\xi}}$ and an arbitrary density $g$ is given by
@@
3 Robust Generalised Bayesian Inference via von Mises-Fisher models
Based on Bissiri et al., (2016); Bissiri and Walker, (2019), the general Bayesian framework achieves belief updating while satisfying coherence and additivity for parameters of interest in the $M$-estimation. In this framework, belief updating is given through a generalised posterior distribution expressed in a functional form. This implies that the generalised posterior provides a probabilistic quantification of the uncertainty associated with the $M$-estimator. Consequently, we apply generalised Bayesian inference to the von Mises-Fisher models using robust divergences.
3.1 Generalised Posterior distributions based on robust divergences
@@ 3.2 Asymptotic Properties
as $n\rightarrow\infty$, where $\phi(\cdot;A)$ is the density function of a p-variate normal distribution with zero mean vector and covariance matrix $A$, and $\pi^{*(d)}(\bm{t}_{n}\mid\bm{x}_{1:n})$ is the generalised posterior density function of the normalized random variable $\bm{t}_{n}=(t_{1},\ldots,t_{p})^{\top}=\sqrt{n}(\bm{\xi}-\hat{\bm{\xi}}_{n}^{(d)})$ given $\bm{x}_{1:n}$.
@@ 3.3 Posterior Computation
In this section, we consider approximate posterior sampling by optimising randomly weighted likelihood functions using the weighted Bayesian bootstrap (WBB) method. Generally, sampling from the posterior distribution is typically performed using MCMC methods, such as Hamiltonian Monte Carlo, the Metropolis-Hastings algorithm, and Gibbs sampling. In directional statistics, numerous studies have developed MCMC for the von Mises-Fisher distribution (Nuñez-antonio and Gutiérrez-peña,, 2005; Pal and Gaskins,, 2022), and a general approach to MCMC simulation on embedded Riemannian manifolds was introduced by Byrne and Girolami, (2013); Habeck et al., (2023). However, implementing these MCMC methods for the robust divergence-based generalised posterior distributions considered in this study is extremely difficult, because the resulting generalised posterior involves special functions and therefore has a highly complicated form. Although tools such as Stan (Carpenter et al.,, 2017) have made MCMC much easier to use in recent years, they cannot be applied in our setting: the DPD and $\gamma$-D posteriors for the von Mises-Fisher models involve Bessel functions of order (p2)/2(p-2)/2, and Stan’s built-in Bessel functions do not support fractional orders. Therefore, Stan cannot be used when the dimension $p$ is odd.
@@ 4 Robustness
Under the quadratic loss function, the DPD Bayes functional $T_{n}^{(\alpha)L}(G)$ and the $\gamma$-D Bayes functional $T_{n}^{(\gamma)L}(G)$ are the DPD posterior mean functional $T_{n}^{(\alpha)e}(G)$ and the $\gamma$-D posterior mean functional $T_{n}^{(\gamma)e}(G)$, respectively. $T_{n}^{(\alpha)e}(G)$ and $T_{n}^{(\gamma)e}(G)$ are given as
@@ 6 Application to real data
In this section, we apply the proposed methods to two real datasets to illustrate their behaviour in real data applications. The R code supporting the results presented in this section is available at https://github.com/nakagawa96/robust-gb-vmf.
@@ 6.1 Wind direction data
We used the wind dataset provided in the circular R package (Agostinelli and Lund,, 2025) to illustrate the basic characteristics of directional data. From Figure 6, while most observations are concentrated around 0 radians, a non-negligible number of observations lie between $\pi/2$ and $\pi$ radians and may be regarded as potential outlying observations relative to the main concentration. Figure 6 illustrates the posterior mean directions of $\bm{\mu}=(\mu_{1},\mu_{2})^{\top}$ estimated from real data using three different posteriors: the ordinary posterior (KL), the DPD posterior, and the $\gamma$-D posterior. Each arrow represents the estimated posterior mean direction on the unit circle, while the surrounding points indicate the sample observations. The tuning parameters $\alpha$ and $\gamma$ are chosen so that ARE is $95\%$, see Appendix A.3 for details.
From Figure 6, we can see that the ordinary posterior mean (KL; $0.29$ radian) is noticeably shifted due to the influence of outliers, whereas the DPD posterior mean (DPD; $0.17$ radian) and $\gamma$-D posterior mean (Gam; $0.16$ radian) remain stable and close to the dominant direction of the data, demonstrating their robustness. The circular dispersion of the posterior samples for $\bm{\mu}$ was smallest for the DPD and $\gamma$-D posteriors, with mean resultant lengths of $0.9993$ and $0.9993$, compared to $0.9989$ for the KL posterior. Consistently, the estimated concentration parameters $\kappa$ were higher for the DPD and $\gamma$-D posteriors ($2.99$ and $3.69$, respectively) than for the ordinary posterior ($1.78$), confirming their robustness to potential outlying observations.
@@ 6.2 Gene expression data
Directional data also arise naturally in bioinformatics (e.g. Eisen et al.,, 1998), particularly in the analysis of gene expression data, in which the von Mises model has been applied for modelling directional structures (e.g. Banerjee et al.,, 2005). We used yeast gene expression data from the NCBI Gene Expression Omnibus (accession GSE12345) (Edgar et al.,, 2002), retrieved via the R package GEOquery (Davis and Meltzer,, 2007), and processed using Biobase from the Bioconductor project (Gentleman et al.,, 2004). To interpret gene expression profiles as directional data, each observation $\bm{z}_{i}=(z_{i1},\ldots,z_{ip})^{\top}$ is normalised to a vector $\bm{x}_{i}=(x_{i1},\ldots,x_{ip})^{\top}$ on the unit sphere by
@@
where $\bar{z}=p^{-1}\sum_{k=1}^{p}z_{ik}$. This normalisation ensures that $|\bm{x}_{i}|_{2}=1$, so that each sample lies on the unit sphere. Under this transformation, the Pearson correlation between $\bm{z}_{i}$ and $\bm{z}_{j})$ can be equivalent to the cosine similarity between $\bm{x}_{i}$ and $\bm{x}_{j}$.
@@
Table 1 presents the angular distance and cosine distance between posterior mean directions obtain from the ordinary, DPD, and $\gamma$-D posteriors ($\alpha=\gamma=0.15$). Figure 8 also provides histograms of the cosine similarities between the data and these posterior mean directions. From the table and figure, we can see that the posterior mean directions are close to one another, suggesting minimal differences in the resulting point estimates. On the other hand, Table 2, which summarises the posterior means of $\kappa$, shows that the ordinary posterior markedly underestimates $\kappa$ due to the strong influence of outliers. By contrast, the DPD and $\gamma$-D posteriors yield considerably more robust estimates.